### v2

# Coexisting phases in the chiral transition within the Linear sigma model with quarks

Abstract
It is believed at present that the chiral transition changes from a smooth crossover to a first-order transition at low temperatures and high densities. Such regime is commonly analyzed using effective models since first principle calculations, as in lattice arrangements, are not feasible. This transition is assumed to be discontinuous, with unstable or metastable intermediate states. However, if multiple charges are simultaneously conserved the system could undergo a continuous change through a coexistence of equilibrium states. This type of transition has multiple manifestations, as in the nuclear liquid-gas transition causing the spinodal fragmentation. The coexistence of phases in the chiral transition is studied here for quark matter assuming the conservation of the isospin composition. Using the Linear sigma model with quarks several remarkable effects are found and discussed.
I INTRODUCTION
The study of matter under extreme conditions of density and temperature has been intensively developed in the last decades with the aim of scrutinizing the different regimes of the strong interaction. Experiments on heavy ion collisions have provided a lot of evidence to understand the high temperature and low matter concentration domain. Complementary information has been obtained from the observational data of compact stars, which corresponds to low temperature and medium-high densities. Some additional physical precisions have been obtained from the lattice simulation of the fundamental theory (LQCD) Alford et al. (1999); Son and Stephanov (2001); Nishida (2004); Ejiri et al. (2010). This method is particularly appropriate to study the vacuum properties and finite temperature effects, although finite density systems are not accessible due to the well known problem of the negative sign. To avoid such failure, different approaches have been proposed, as for instance the use of an imaginary baryonic chemical potential $\mu_{B}$, or the introduction of an isospin chemical potential $\mu_{I}$ Alford et al. (1999), which takes into account the conservation of the flavor degree of freedom. Furthermore, theoretical investigations using effective models have tried to compile the phenomenology within a unified description. As an outcome of these common efforts, a picture of a complex phase diagram has emerged, with several phase transitions, collective states, etc.
Special attention was paid to the deconfinement phase transition Aryal et al. (2020); Prakash et al. (2024) since its realization could have clear manifestations in heavy ion collision events or, as pointed out more recently, it could leave its imprint in the gravitational waves coming from the collapse of companion neutron stars.
The fluctuations of conserved charges, such as baryon number or electric charge, have been proposed as an efficient signal of the transition and consequently have been focused by the LQCD . They have been related to the corresponding thermodynamical susceptibilities by the fluctuation dissipation theorem and have been used to reconstruct the equation of state.
The chiral phase transition has also been an object of theoretical investigation. As in the case of the deconfinement, the general consensus is that its character changes from first order at low temperatures to second order at higher ones. In the case that non zero quark mass is considered, the second order transition becomes a soft crossover. As a consequence, the line of transition points in the $(\mu_{B},T)$ plane exhibits a critical endpoint (CEP) separating both regimes.
Systems at zero baryon density ($\mu_{B}=0$) but with an isospin imbalance ($\mu_{I}\neq 0$) have been hypothesized to be in a pion condensed phase before the chiral symmetry is restored and $\mu_{I}>m_{\pi}$ Son and Stephanov (2001). This is an interesting assumption that relies on the identification of $\mu_{I}$ as a gauge field. However, it was appropriately warned that a system with $\mu_{I}>\mu_{B}$ will be physically unstable by weak decay Son and Stephanov (2001).
General features of the chiral transition in quark matter with isospin imbalance have been investigated in Toublan and Kogut (2003); Frank et al. (2003); Barducci et al. (2004, 2003); Loewe and Villavicencio (2005); He et al. (2005); He and Zhuang (2005); Zhang and Liu (2007); Mukherjee et al. (2007); Kovacs and Szep (2008); Mu et al. (2010); Abuki (2013); Xia et al. (2013); Stiele et al. (2014); Liu et al. (2016); Avancini et al. (2019); Lopes et al. (2021); Adhikari and Andersen (2020); Costa et al. (2020); Chu et al. (2023); Ayala et al. (2023, 2025). The coherence of the theoretical predictions with the LQCD results at zero baryon number has been particularly analyzed in Loewe and Villavicencio (2005); Avancini et al. (2019); Lopes et al. (2021); Adhikari and Andersen (2020); Ayala et al. (2025).
In this paper, instead, quark matter at finite baryon density and with isospin asymmetry is considered. For this purpose an effective model, the Linear Sigma Model with quarks (LSMq), is adopted and the weak interaction is neglected as in Son and Stephanov (2001). Thus, the $u$ and $d$ quark flavors can be regarded as stable degrees of freedom, and an additional meson $\zeta$ with isospin structure is introduced and treated on the same foot as the standard $\sigma,\pi$ ones. Several attempts to enlarge the meson family within the LSMq have been made Parganlija et al. (2010, 2013); Kovács et al. (2016), particularly the role of the scalar, isovector meson has been discussed from long time ago Gasiorowicz and Geffen (1969); Metzger et al. (1994); Lenaghan et al. (2000); Schaefer and Wagner (2009); Abuki (2013); Ayala et al. (2020). In order to stress the results on thermodynamical instabilities, the simplicity of the theoretical formulation is preferred here. Hence, only the meson fields $\sigma,\,\pi,\,\zeta$ are considered, and a more sophisticated approach is left for further development.
The remaining of this work is organized as follows, the mean field approach to the LSMq, also known as Quark Meson model, is presented in the next section. The results for the thermodynamics of the system is discussed in Sec. III, and the final conclusions are drawn in Sec. IV.
II THE MODEL
The LSMq is an effective model of the strong interaction, adequate for the low energy regime. It has been used extensively due to its versatility and simplicity. The concise Lagrangian density used in this work is Ayala et al. (2020)
${\cal L}$ $=$ $\,\bar{\Psi}\left(i\not\!\partial+g\,\Phi\right)\Psi+\frac{1}{2}\left(\partial_{\mu}\sigma\partial^{\mu}\sigma+\partial_{\mu}\bm{\pi}\cdot\partial^{\mu}\bm{\pi}+\partial_{\mu}\bm{\zeta}\cdot\partial^{\mu}\bm{\zeta}\right)+\frac{1}{4}C_{0}\text{Tr}\left(\Phi^{\dagger}\Phi\right)$ (1)
$-\frac{1}{16}C_{3}\left[\text{Tr}\left(\Phi^{\dagger}\Phi\right)\right]^{2}-\frac{1}{8}C_{2}\text{Tr}\left(\Phi^{\dagger}\Phi\right)^{2}-h\,\sigma$
with the quark bi-spinor $\Psi=\left(\psi_{u}\;\psi_{d}\right)^{t}$, the iso-multiplets scalar $\bm{\zeta}$, and pseudo-scalar $\bm{\pi}$, and $\Phi=\sigma+i\,\gamma_{5}\bm{\pi}\cdot\bm{\tau}+\bm{\zeta}\cdot\bm{\tau}$. Flavor degeneracy in the quark masses is assumed and a symmetry breaking term, linear in $\sigma$ has been included.
The scalar meson fields can be decomposed as the sum of a thermal expectation value and its fluctuation $\sigma=s+\delta\sigma$, $\zeta_{a}=z\,\delta_{a3}+\delta\zeta_{a}$. Neglecting the fluctuations leads one to the mean field approach ($MFA$), while its explicit consideration can be arranged as higher order corrections Dolan and Jackiw (1974); Petropoulos (1999); Cornwall et al. (1974). Here I adopt the one-loop approach, as in Ayala et al. (2025), where mesons behave as free particles with effective masses. Hence the grand potential per unit volume of uniform quark matter can be written as
$\omega(T,\mu)$ $=$ $\omega_{\text{vac}}-\frac{N_{c}}{\beta\pi^{2}}\sum_{j=u,d}\int_{0}^{\infty}dp\,p^{2}\left[\text{log}\left(1+e^{-\beta(E_{p}-\mu_{j})}\right)+\text{log}\left(1+e^{-\beta(E_{p}+\mu_{j})}\right)\right]$ (2)
$+\frac{1}{2\beta\pi^{2}}\sum_{\alpha=\sigma,\pi_{a},\zeta_{a}}\int_{0}^{\infty}dp\,p^{2}\,\text{log}\left(1-e^{-\beta\,E_{\alpha}}\right)$
The vacuum term is divergent, but a finite contribution can be extracted by dimensional regularization followed by an appropriate subtraction scheme
$\omega_{\text{vac}}=\frac{N_{c}}{8\pi^{2}}\sum_{j=u,d}\left[m_{j}^{4}\,\text{log}\left(\frac{m_{j}}{m_{0}}\right)+\frac{m_{0}^{4}-m_{j}^{4}}{4}\right]-\frac{1}{32\pi^{2}}\sum_{\alpha=\sigma,\pi_{a},\zeta_{a}}\left[M_{\alpha}^{4}\,\text{log}\left(\frac{M_{\alpha}}{M_{0\alpha}}\right)+\frac{M_{0\alpha}^{4}-M_{\alpha}^{4}}{4}\right]$ (3)
The notation $m_{j}=g(s+I_{j}\,z)$ has been introduced for the effective quark masses, where $I_{j}=1\,(-1)$ for j=u(d)j=u\,(d). The subindex 0 in the masses indicates the physical values.
Within the formalism of Cornwall et al. (1974); Amelino-Camelia and Pi (1993), the effective potential is a functional of the mean values s,zs,\,z and of the full propagators of the mesons. Hence the definitions of s,zs,\,z are given,as usual, by the extremum conditions
$0=\frac{\partial\omega}{\partial s},\;\;\;\;\;\;0=\frac{\partial\omega}{\partial z}$
Giving
$0=(-C_{0}+C_{3}s^{2}+C_{4}z^{2})\,s-h-\frac{gN_{c}}{\pi^{2}}\sum_{j=u,d}m_{j}\left[\frac{m_{j}^{2}}{2}\,\text{log}\left(\frac{m_{j}}{m_{0}}\right)-\int_{0}^{\infty}\frac{dp\;p^{2}}{E_{j}}\left(n_{Fj}+\bar{n}_{Fj}\right)\right],$ (4)
$0=(-C_{0}+C_{3}z^{2}+C_{4}s^{2})\,z-\frac{gN_{c}}{\pi^{2}}\sum_{j=u,d}I_{j}\,m_{j}\left[\frac{m_{j}^{2}}{2}\,\text{log}\left(\frac{m_{j}}{m_{0}}\right)-\int_{0}^{\infty}\frac{dp\;p^{2}}{E_{j}}\left(n_{Fj}+\bar{n}_{Fj}\right)\right],$ (5)
respectively. In addition, there is a condition for each of the mesons, which in the present approach reduces to $M_{\sigma}^{2}=-C_{0}+3C_{3}\,s^{2}+(C_{3}+2C_{2})\,z^{2}$, $M_{\pi\,a}^{2}=-C_{0}+C_{3}(s^{2}+z^{2})$, $M_{\zeta\,a}^{2}=-C_{0}+C_{3}\,n_{a}\,z^{2}+(C_{3}+2C_{2})\,s^{2}$ and $n_{a}=1\,(3)$ for a=1,2(3)a=1,2\,(3).
In Eqs.(4, 5) the equilibrium Fermi distribution functions $n_{Fj}(T,\mu_{j})$ for quarks and $\bar{n}_{Fj}(T,\mu_{j})$ for antiquarks have been introduced.
The entropy of the system is given by the thermodynamical definition ${\cal S}=S/V=-\partial\,\omega/\partial T$ as
${\cal S}(T,\mu)$ $=$ $-\frac{\omega-\omega_{\text{vac}}}{T}+\frac{N_{c}}{\pi^{2}T}\sum_{j=u,d}\int_{0}^{\infty}dp\,p^{2}\left[(E_{j}-\mu_{j})\,n_{Fj}+(E_{j}+\mu_{j})\,\bar{n}_{Fj}\right]$ (6)
$+\frac{1}{2\pi^{2}T}\sum_{\alpha=\sigma,\pi_{a},\zeta_{a}}\int_{0}^{\infty}dp\,p^{2}\,E_{\alpha}\,n_{B\alpha}.$
In the last equation the equilibrium Bose distribution function $n_{B}(T)$ is used. Finally, the energy density is evaluated by the Legendre transform ${\cal E}=\omega+T\,{\cal S}+\sum_{j=u,d}\mu_{j}\,n_{j}$, using the quark number density
$n_{j}=\frac{N_{c}}{\pi^{2}}\int_{0}^{\infty}dp\,p^{2}\left(n_{Fj}-\bar{n}_{Fj}\right).$
In the following the evolution of matter at fixed isospin fraction $x=(n_{d}-n_{u})/(n_{u}+n_{d})$ is considered. Some characteristic quantities of the thermodynamical evolution of the system are the isothermal speed of sound $v=-\left(\partial\omega/\partial{\cal E}\right)_{T}$ and the second order susceptibilities $\chi$ Gavai and Gupta (2001); Ferroni and Koch (2011); Ratti et al. (2007); Sasaki et al. (2007a, b, 2008); Stokic et al. (2009); Skokov et al. (2010a); Almasi et al. (2017); Skokov et al. (2010a, b, 2011). The former has been demonstrated to be useful to indicate the emergence of new degrees of freedom. Meanwhile the susceptibilities manifest the dynamical fluctuations characterizing the bulk (semi-classical) behavior of the system Asakawa et al. (2000); Gavai and Gupta (2006); Datta et al. (2017). Fluctuations are closely related to phase transitions, in particular those related to conserved charges. For such reason the susceptibilities associated with conserved charges have been focused within the LQCD Ratti et al. (2007); Gavai and Gupta (2006); Datta et al. (2017).
In the following the susceptibilities associated with the quark numbers
$\chi_{B}=\left(\frac{\partial n_{B}}{\partial\mu_{B}}\right)_{\mu_{3}},\;\;\chi_{3}=\left(\frac{\partial n_{3}}{\partial\mu_{3}}\right)_{\mu_{B}},$
will be considered. Here the densities for the baryon number $n_{B}=(n_{d}+n_{u})/3$, and the isospin number $n_{3}=n_{d}-n_{u}$ have been introduced, together with the chemical potentials for the baryon $\mu_{B}=3(\mu_{d}+\mu_{u})/2$ and the isospin $\mu_{3}=\mu_{d}-\mu_{u}$ charges.
III RESULTS AND DISCUSSION
The parameters shown in Eq. (1) depend on the physical masses of the mesons and the pion decay constant as
$C_{0}=\frac{M_{\sigma 0}^{2}-3M_{\pi 0}^{2}}{2},\;C_{2}=\frac{M_{\zeta 0}^{2}-3M_{\pi 0}^{2}}{2f_{\pi}^{2}},C_{3}=\frac{M_{\sigma 0}^{2}-M_{\pi 0}^{2}}{2f_{\pi}^{2}},$
and the values $M_{\sigma 0}=500$ MeV, $M_{\pi 0}=138$ MeV, and $M_{\zeta 0}=984$ MeV have been used. In particular the mass of the $\zeta$ meson has been identified with that of the $a_{0}$(980) since this is the lightest manifestation in the hadronic sector of an scalar isoscalar. The coupling constant $g$ is related to the quark constituent mass through $m_{q0}=g\,f_{\pi}$ and will be defined in the next step.
The model is used in the following to analyze the chiral phase transition in homogeneous quark matter. The relation between chiral breakdown and deconfinement is not clear yet, and for the sake of simplicity it is not treated here. However, this subject has been extensively studied, specially with the aid of a phenomenological Polyakov potential. Under the conditions considered in this work the isospin chemical potential verifies $\mu_{3}<M_{\pi 0}$, for this reason the pion condensation is not taken into account.
As finite density matter is considered the approach turns inadequate since the pion mass becomes imaginary. As for example, in flavor symmetric matter it is found that $M_{\pi}^{2}<0$ as soon as $s/f_{\pi}<0.9$. In consequence the contributions of all the mesons, as in Eqs.(2), (3), (6), are disregarded in the following but keeping the one-loop approach for quarks. A similar treatment is usual in relativistic hadronic models, and it is commonly applied to study phase transitions as the nuclear liquid-gas, or deconfinement in high density matter. The MFA is also the standard approach in quark matter calculations using the effective Nambu-Jona Lasinio model Sasaki et al. (2007a, b, 2008), and it has been used in the case of the LSMq model too Stokic et al. (2009).
At present it is believed that the chiral transition is a continuous crossover for low densities and becomes first order for sufficiently high values of $\mu_{B}$. There is no clear definition of the transition temperature in a crossover, but to show a characteristic temperature we adopt the value for the inflection point of the chiral order parameter Pawlowski and Rennecke (2014), i.e. where the minimum of $s^{2}+z^{2}$ occurs. The critical temperature $T_{0}$ is found to be a decreasing function of $g$ and low values of $m_{q0}$ favors the compatibility with LQCD estimations Ejiri et al. (2010). However, in the present approach it is found that for the lower $m_{q0}$ the expected change of regime for high $\mu_{B}$ is lost. Therefore an intermediate value is adopted $m_{q0}=250$ MeV, which yields the phase diagram shown in Fig.1 for several values of $x$. The value obtained $T_{0}\simeq 120$ MeV is out the range 160180160-180 MeV predicted by LQCD, but a better fit could be obtained by including higher order corrections to the present scheme Petropoulos (1999); Ayala et al. (2020). In regard of the values selected for the flavor asymmetry, $x=0$ corresponds to the flavor symmetric case usually discussed in finite density calculations, $x=1/3$ represents electric charge neutral matter similar to that found in self-bound quark stars, and finally $x=2/3$ is an extrapolation to the negatively charged case.
As expected a change of regime is found with a CEP in between. For $T=0$ and $\mu_{B}=0$ all the curves coincide, while at intermediate values of the chemical potential the pseudo-critical temperature decreases with $x$.
The first order phase transition is generated by instabilities in the equation of state, since the thermodynamic potential $\omega$ does not vary monotonically with the chemical potentials. Depending on the value of the surface tension, the transition occurs as a discontinuous change or as a continuous passage with an intermediate coexistence region. The last case is feasible because there are multiple charges conserved simultaneously Glendenning (1992); Heiselberg and Hjorth-Jensen (2000), and both coexisting phases correspond to the regime of broken chiral symmetry. Assuming a low surface tension the system evolves by following the Gibbs construction, and a parameter $\lambda$ is introduced which indicates the relative abundance of the coexisting phases, named $a$ and $b$ in the following. Hence the equilibrium conditions read
$T_{a}=T_{b},\;\;\mu_{B}^{a}=\mu_{B}^{b},\;\;\mu_{3}^{a}=\mu_{3}^{b},$ (7)
$P_{a}(T_{a},\mu_{B}^{a},\mu_{3}^{a})=P_{b}(T_{b},\mu_{B}^{b},\mu_{3}^{b}),$ (8)
while for every additive thermodynamical function one has
$n_{B}=\lambda\,n_{B}^{a}+(1-\lambda)\,n_{B}^{b},\;{\cal S}=\lambda\,{\cal S}^{a}+(1-\lambda)\,{\cal S}^{b},\;{\cal E}=\lambda\,{\cal E}^{a}+(1-\lambda)\,{\cal E}^{b},\;\text{etc.}$
$0\leq\lambda\leq 1$. To evaluate the derivatives of the thermodynamical potential correctly, the dependence of the parameter $\lambda(T,\mu)$ must be taken into account.
Due to the fact that the thermodynamical potential $\omega(T,\mu_{B},\mu_{3})=-P$ is non monotonous, the states $a$ and $b$ have common values of temperature and chemical potentials and yet different mean field values $s,z$ and flavor compositions. It must be stressed that for $x=0$ there is only one conserved charge, since in such case the flavors u,du,\,d are indistinguishable. Therefore the Gibbs construction does not apply. For other global asymmetries $x$, pairs of states with different isospin compositions $x_{a},\;x_{b}$ are associated under the constraint
$x=\frac{\lambda\,x_{a}\,n_{a}+(1-\lambda)\,x_{b}\,n_{b}}{\lambda\,n_{a}+(1-\lambda)\,n_{b}}.$ (9)
For a fixed temperature these states form a continuum with extreme values $x_{L}$ and $x_{H}$. For a given pressure the collection of all these points, corresponding to the full range of temperatures allowed, constitutes the Equilibrium Coexistence Region (CER) region. To illustrate this subject the ECR corresponding to $P=43.5$ MeV fm-3 is shown in Fig. 2. It extends up to a maximum temperature $T=9.2$ MeV, where only states with $x_{a}\simeq x_{b}\simeq 0$ are involved. As a general trend it is found that the extension of this region reduces when i) the pressure increases, or ii) $g$ decreases, until it completely disappears.
An interesting quantity which manifests some features of the equation of state is the speed of sound $v_{S}$. It has been proposed that the value of $v_{S}$ can be inferred from experimental data of ultrarelativistic heavy ion collisions Gardim et al. (2020); Mu et al. (2025). It also enters in the definition of the tidal deformability of binary star systems and non-radial oscillations of compact stars. For this reason it can indirectly be estimated by the analysis of the gravitational waves, whose detection is under permanent attention due to the evolution of methods and technologies. So, there is a strong motivation for the recent studies of the speed of sound in quark matter using effective models Ayala et al. (2023, 2025); Ghosh et al. (2006); He et al. (2005); He and Zhuang (2005); Liu et al. (2016); Hernández-Ortiz et al. (2024). Much of them explore the limit of $\mu_{B}\rightarrow 0$ in order to contrast with LQCD results, but in Hernández-Ortiz et al. (2024) the relation between the speed of sound and the CEP is studied.
In Fig.3 the isothermal speed of sound is shown for a range of quark densities and temperatures corresponding to the first-order transition and two flavor compositions x=1/3(2/3)x=1/3\,(2/3) in the upper (lower) panel. For $x=1/3$ and $T=20$ MeV the temperature is above the CEP and $v_{S}$ has a non-monotonic trend with its minimum located approximately at the transition point (filled circle). For the lower temperatures the system experiences the coexistence of phases, which reflects in a discontinuous behavior with a sudden drop of $v_{S}$. Along this passage the speed of sound has a mild variation with the particle number, and it is approximately independent of the temperature. Furthermore, the shrinking of the ECR region as $T$ grows is evident. A similar description applies to $x=2/3$, but in such case only the isotherms corresponding to $T=0,\text{and}\;5$ MeV undergo a first order transition.
To end this section, the results for the susceptibilities associated with the conserved charges will be presented. The effects of the chiral transition on the number susceptibilities have been discussed previously within the NJL model Sasaki et al. (2007a, b, 2008), and also with the LSMq Stokic et al. (2009); Skokov et al. (2010a, b, 2011); Almasi et al. (2017), and particularly the presence of a spinodal out of thermodynamical equilibrium has been considered Sasaki et al. (2007b, 2008).
Along the last figures the susceptibilities are shown as functions of the quark density for several temperatures in the range of the first order transition. For $\chi_{B}$ (Fig.4) strong oscillations are found, concentrated at the borders of the binodal. For higher densities all the curves decay and reach the asymptotic regime quickly. It must be pointed out that within a Maxwell construction this dramatic behavior is lost since most of this range of densities is not accessible for the system.
For $x=0$ (Fig.4a) results including the unstable region of the equation of state are shown, the peaks and drops are truly divergences of the same type as those reported and discussed in Sasaki et al. (2007b, 2008) for calculations using the NJL model.
For $x>0$ the figures are more complex because, in contrast to $x=0$, the system experiences a coexistence of phases. This effect is clear for $x=1/3$ (Fig.4b) where the steep peaks and drops (dashed lines) are replaced by smooth curves with finite discontinuities at the end points. In between $\chi_{B}$ remains positive and increasing with the density, a fact that is emphasized with growing $T$.
For the higher $x=2/3$ a qualitative change happens. For $T=10$ MeV the system undergoes a continuous crossover and $\chi_{B}$ shows a regular behavior. The cases $T=0,\text{and}\;5$ MeV, show noticeable changes since even within the ECR $\chi_{B}$ has large but finite discontinuities. However it must be bear in mind that this is an extrapolation, since the effect of the weak forces will not be negligible for $x=2/3$.
The susceptibility associated with isospin is completely regular for $x=0$ (Fig.5a) with an evident enhancement for non-zero temperatures. For $x=1/3$ (Fig.5b) the unstable region is marked by peaks and drops in $\chi_{3}$, which are replaced by a monotonous trend within the ECR. Finally, for $x=2/3$ (Fig.5c) the significant deviations from the null behavior only occurs at the high-density limit of the binodal and are distinguished by a divergence for $T=5$ MeV. However the same remarks as for $\chi_{B}$, $x=2/3$ apply here.
IV SUMMARY AND CONCLUSIONS
The behavior of the chiral phase transition under asymmetric flavor composition of the quark matter has received great attention in recent years. One of the causes is that LQCD calculations have been available, and this is an opportunity for effective models to improve their schemes by contrasting predictions with well established results. Since LQCD is efficient in the limit of zero matter density, much of the recent theoretical efforts have been restricted to such regime Son and Stephanov (2001); Stiele et al. (2014); Avancini et al. (2019); Lopes et al. (2021); Adhikari and Andersen (2020); Ayala et al. (2023, 2025). However, interesting effects have been pointed out in regard of the chiral transition at non-zero baryonic density. In particular the effects of fluctuations near the non-equilibrium region of the first order phase transition have been studied Sasaki et al. (2007a, b, 2008); Stokic et al. (2009); Skokov et al. (2010a, b, 2011); Almasi et al. (2017). By definition this domain can not be analyzed by the equilibrium thermodynamics. However, if multiple charges are conserved and a low surface tension is assumed, the system could undergo a continuous transition through the coexistence of equilibrium states Glendenning (1992); Heiselberg and Hjorth-Jensen (2000). These hypothesis have been developed in this paper in the context of the LSMq with two flavors. The scalar isoscalar meson $\zeta$ has been included as an emergent of the isospin imbalance at finite density. The one loop approximation has been proposed, but calculations have been finally performed in the MFA. The effects of higher order corrections is object of present calculations. Neither pion condensation nor quark superconductivity have been considered since the chemical potentials $\mu_{B},\;\mu_{3}$ are below the corresponding thresholds for the range of this calculations.
The phase diagram obtained is in qualitative agreement with present knowledge, a smooth crossover at high temperatures ends at a CEP and is replaced by a first order transition at low temperatures. Although the value obtained for the critical temperature is too low, it is expected that higher order corrections will improve this result Petropoulos (1999); Ayala et al. (2020).
Several important effects on the bulk properties of quark matter with isospin imbalance have been found and discussed. The speed of sound at finite isospin and temperatures above the CEP, have a non-monotonic dependence on the particle density exhibiting a minimum corresponding to the crossover transition. Within the coexistence region $v_{S}$ experiences a noticeable decrease with finite discontinuities at the borders. A mild variation with density and temperature is predicted for this domain.
The susceptibility associated with the baryon number $\chi_{B}$, remains bounded for low asymmetry parameter $x>0$, in contrast to calculations using the unstable sector of the equation of state. As $x$ grows large but finite discontinuities appear at the high density border of ECR, but the inclusion of the electromagnetic field could have important effects in such situations. A similar comment applies to the susceptibility $\chi_{3}$, corresponding to the isospin asymmetry.
Acknowledgements
This work has been partially supported by CONICET, Argentina under the project 11220200102081CO, and by UNLP, Argentina.